You are a manager at Glass Inc.—a mirror and window supplier. Recently, you conducted a study of the production process for your single-side encapsulated window. The results from the study are summarized in the table below, and are based on the 8 units of capital currently available at your plant. Workers are paid $60 per unit, per unit capital costs are $20, and your encapsulated windows sell for $12 each. Given this information, optimize your human resource and production decisions. Do you anticipate earning a profit or a loss?explain carefully
labour output
0 0
1 10
2 30
3 60
4 80
5 90
6 95
7 95
8 90
9 80
10 60
11 30
1
Expert's answer
2019-10-31T10:36:04-0400

The mirror and window supplier is a profit-maximizing firm that will hire labor up to the point where the value of its marginal product equals the wage rate of $60. There'll are currently 8 units of capital available at the plant.

Firm’s demand for labor is the only decision that the firm is conscious about. Given the wage rate of $60 per unit of labor, the firm will continue to hire labor till the value of marginal product of labor reaches $60.


labour output MP AP

0 0 0 0

1 10. 10-0=10 10/1=10

2 30. 30-10=20. 30/2=15

3 60. 60-30=30. 60/3=20

4 80. 80-60=20. 80/4=20

5 90. 90-80=10. 90/5=14

6 95. 95-90=5. 95/6=15,8

7 95. 95-95=0. 95/7=13,5

8 90. 90-95=-5. 90/8=11,25

9 80. 80-90=-10. 80/9=8,89

10 60. 60-80=-20. 60/10=6

11 30. 30-60=-30. 30/11=2,9


Thus maximizing productivity and return on resources Observed when hiring 3 and 4 employees Profit maximization will be observed if 4 employees are hired:

80 * 12-60 * 4-20 * 8 = $ 560